OpenStudy (maggy14145):

The system of equations graphed below will have _____ positive solutions. Enter the number that correctly fills in the blank in the previous sentence.

jimthompson5910 (jim_thompson5910):

A solution is visually represented by the point of intersection between the two curves There are 6 intersection points, so 6 solutions. However, ALL of the intersection points have x coordinates that are negative. Notice how everything is to the left of the vertical y axis. This means that they are considered `negative solutions` and they are NOT positive solutions So there are 0 positive solutions

jimthompson5910 (jim_thompson5910):

The two curves would have to intersect on the right side of the y axis for a positive solution to occur.

OpenStudy (maggy14145):

wow ok thank you. I have a similar question using the same graph and I already think my answer is correct but I just want to make sure with you if thats ok with you :)

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (maggy14145):

The system of equations graphed below appears to have _____ solutions. Enter your answer as the number that correctly fills in the blank in the previous sentence.

OpenStudy (maggy14145):

so the answer for this would be 6 correct?

OpenStudy (maggy14145):

using the same graph of course

jimthompson5910 (jim_thompson5910):

yes if you drop the `positive` term from the sentence, then there are 6 solutions overall

OpenStudy (maggy14145):

GOD BLESS YOUR SOUL THANK YOU SO MUCH

OpenStudy (maggy14145):

would you mind answering just one more question..? I'm very confused..

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (maggy14145):

All four functions below are quadratic functions. Which function has the greatest maximum value?

OpenStudy (maggy14145):

y= -(x-10)^2-4

jimthompson5910 (jim_thompson5910):

it only gives you one function?

OpenStudy (maggy14145):

no sorry I'm trying to attach the other options give me like 1 min

OpenStudy (maggy14145):

y= -x^2-6x-17

OpenStudy (maggy14145):

those are the 4 options

jimthompson5910 (jim_thompson5910):

oh gotcha

jimthompson5910 (jim_thompson5910):

y= -(x-10)^2-4 is in the form y = a(x-h)^2 + k where a = -1 h = 10 k = -4 do you agree?

OpenStudy (maggy14145):

correct

jimthompson5910 (jim_thompson5910):

the value of k will determine either the min value of y or the max value of y if a > 0, then k will be the min if a < 0, then k will be the max a = -1 means a < 0, so k = -4 is the max of y= -(x-10)^2-4 the largest that y can get is -4

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (maggy14145):

yes. slowly but surely following

jimthompson5910 (jim_thompson5910):

So write this down: the max of choice A is -4. We'll come back to this later

jimthompson5910 (jim_thompson5910):

looking at this graph http://assets.openstudy.com/updates/attachments/57a512bde4b035cb2d6aef9c-maggy14145-1470436674415-stup.png where is the peak located? What is the highest point?

OpenStudy (maggy14145):

16, correct..?

jimthompson5910 (jim_thompson5910):

it's located at (1,16) (x,y) coordinates help you figure out location the y value is 16, so the max is 16

jimthompson5910 (jim_thompson5910):

choice A) The max is -4 choice B) The max is 16 so far, B is winning the race

OpenStudy (maggy14145):

ok I see that

jimthompson5910 (jim_thompson5910):

Look at choice C http://assets.openstudy.com/updates/attachments/57a512bde4b035cb2d6aef9c-maggy14145-1470436696490-stup2.png what's the largest y value shown?

OpenStudy (maggy14145):

19

OpenStudy (maggy14145):

so the answer to this question would be the 3rd option since the last option's y-value is -17, correct?

OpenStudy (maggy14145):

and 19 is considered the highest y-value out of options

jimthompson5910 (jim_thompson5910):

`since the last option's y-value is -17, correct?` no -17 is not the max for choice D

jimthompson5910 (jim_thompson5910):

y= -x^2-6x-17 is the same as y= -1x^2-6x-17 y= -1x^2-6x-17 is in the form y = ax^2+bx+c where... a = -1 b = -6 c = -17 The values of 'a' and 'b' will be plugged into `x = -b/(2a)` what do you get when you do this?

OpenStudy (maggy14145):

-3

jimthompson5910 (jim_thompson5910):

now plug x = -3 back into y= -x^2-6x-17 to get y= -x^2-6x-17 y= -(-3)^2-6(-3)-17 ... replace every x with -3 y = -(9) - 6(-3) - 17 y = -9 + 18 - 17 y = ???

OpenStudy (maggy14145):

-8

OpenStudy (maggy14145):

so the correct answer would be the third option since the rest of the maximum values of the other options were less than the 3rd option right?

jimthompson5910 (jim_thompson5910):

so the max of choice D is -8

OpenStudy (maggy14145):

yes correct

jimthompson5910 (jim_thompson5910):

now I should point out that the max of choice C is NOT 19. It's actually something larger. However, we don't need to find the actual max value. We have this summary Max of choice A is -4 Max of choice B is 16 Max of choice C is some value larger than 19 Max of choice D is -8 so even without knowing the actual max of choice C, we can see that it is the largest max

OpenStudy (maggy14145):

Thank you so much for helping me! I really appreciate it :)

jimthompson5910 (jim_thompson5910):

Glad to be of help